# The Role of Learning Algorithms in Collective Action

### v1

@@ Abstract
Collective action in Machine Learning is the study of the control that a coordinated group can have over machine learning algorithms. While previous research has concentrated on assessing the impact of collectives against Bayes optimal classifiers, this perspective is limited in that classifiers seldom achieve Bayes optimality, and are influenced by the choice of learning algorithms along with their inherent biases. In this work, we initiate the study of how the choice of the learning algorithm plays a role in the success of a collective in practical settings. Specifically, we focus on distributionally robust algorithms (DRO), popular for improving a worst group error, and on the ubiquitous stochastic gradient descent (SGD), due to its inductive bias for “simpler” functions. Our empirical results, supported by a theoretical foundation, show that the effective size and success of the collective are highly dependent on properties of the learning algorithm. This highlights the necessity of taking the learning algorithm into account when studying the impact of collective action in machine learning.
{NoHyper}footnotetext: Email: odov@tue.mpg.de, jake.fawkes@stats.ox.ac.uk, ssamadi@tuebingen.mpg.de, amsa@di.ku.dk{NoHyper}**footnotetext: Equal contribution
@@ 1 Introduction
Hardt et al. (2023) initiated the formal study into the success of different collective action strategies. Their work provided theoretical analysis based on the fractional size of the collective and on the properties of the signal with regards to the original data distribution. However, their results only hold for Bayes optimal classifiers which do not immediately adapt to peculiarities of practical learning algorithms. In this work, we investigate two types of commonly used learning algorithms that exhibit distinct properties: (1) Distributional robustness and (2) Simplicity bias. We show how these properties lead to significantly different levels of collective success that are unexplained by prior work.
When the data is comprised of multiple sub-populations, algorithms that optimise for average performance often perform poorly in minority sub-populations (Meinshausen and Bühlmann, 2015). A fairness or safety-focused firm will want to ensure that their trained learning model performs well uniformly on all sub-populations, as opposed to being on-average good. Distributionally Robust Optimisation (DRO) is a family of algorithms designed to maximize this per-group accuracy (Hashimoto et al., 2018). We show that as a consequence, a small collective achieves much higher success when the training algorithm performs DRO, compared to standard empirical risk minimization (ERM). Conversely, a large collective in the same settings achieves lower success, contradicting the expectation set by previous work regarding the effectiveness of large collectives.
Second, most machine learning algorithms used today are based on some form of gradient descent (GD). Such algorithms, including the popular Stochastic Gradient Descent (SGD), Adam, and RMSProp, exhibit a preference for learning functions that are “simpler". This inductive bias of GD algorithms is popularly referred to as simplicity bias (Kalimeris et al., 2019; Shah et al., 2020). In practice, this preference results in the model “overlooking” certain complex features. We demonstrate that these overlooked features can be leveraged by a collective to design a strategy that will gain higher success compared to what is possible on a Bayes optimal classifier.
@@ 2.1 Collective Action
While there are several problems in the domain of algorithmic collective action defined in Hardt et al. (2023), in this work we will focus on the problem of collectively planting a signal with a features-label strategy. We believe other important problems like erasing signals can also benefit from the techniques and observations of this paper but we leave the detailed study of such settings to future work.
@@
The collective modifies their data by planting a signal $\left(x,y\right)\rightarrow\left(g\left(x\right),y^{*}\right)$${\mathcal{P}}^{*}$ defines the distribution of $\left({g\left({x}\right),y^{*}}\right)$ where $x\sim{\mathcal{P}}_{0}$. This has been defined as the feature-label strategy in prior work. The signal $g$ also defines the signal set ${\mathcal{X}}^{*}=\left\{g\left(x\right)|x\in{\mathcal{X}}\right\}$. For any distribution ${\mathcal{P}}$ over $X{\times}Y$, the Bayes optimal classifier on ${\mathcal{P}}$, which we denote as $f_{\mathcal{P}}$, is defined as
@@
Using the above four properties they derive the following lower bound on the success of the collective.
@@ 2.2 Distributionally Robust Optimisation
where ${\mathcal{Q}}_{p}$ is an uncertainty set of distributions close to $p$ over which we want to control the risk, $\ell$ is the loss function, and $g_{\theta}$ is a function with parameters $\theta$. There are many possible definitions of the uncertainty set that the algorithm is controlling for. One possible choice is an $f$-divergence [^1] A definition of $f$-divergence is provided in definition 2. ball (Ali and Silvey, 1966; Csiszár, 1967).
@@ 3 Effective Size and Validation Control
The most intuitive parameter to predict the success of collective action is the fractional size of the collective $\alpha$; a larger collective will attain greater success. However, DRO algorithms assign different weights to different samples, rendering $\alpha$ inappropriate for predicting success. Instead, we introduce a correction to the collective size under a weighted distribution that we denote the effective size $\alpha_{\text{eff}}$, and show that DRO algorithms yield $\alpha_{\text{eff}}>\alpha$.
@@ 3.1 Effective Collective Size
Note that if $w\left({x}\right){=}1$ for all $x$, then $\alpha_{\text{eff}}{=}\alpha$. Yet, since $\alpha_{\text{eff}}$ is the collective size under the weighted distribution, then bounding success for this algorithm, akin to Theorem 1, requires adding a corrective term $c$ for the non-weighted distribution, giving
@@
where $\Delta_{x}^{w}=\max_{y\in Y}\left({\mathcal{P}}^{\left(w\right)}\left(y\mid x\right)-{\mathcal{P}}^{\left(w\right)}\left(y^{*}\mid x\right)\right)$ and $\Delta_{x}$ is defined the same but for ${\mathcal{P}}_{0}$. Formal proof and definitions are in the appendix under Proposition 7. Under certain circumstances we can have $c\leq 0$, implying a collective success greater than that guaranteed by Theorem 1 for $\alpha=\alpha_{\text{eff}}$. We provide such an example where $c\leq 0$ and $\alpha_{\text{eff}}\geq\alpha$ in Appendix B.1.
We now provide theoretical results, showing that DRO algorithms can increase the effective collective size $S\left(\alpha_{\text{eff}}\right)$. Firstly, for any algorithm that targets a DRO defined by an $f$-divergence, we can say the following:
Proposition 2.
@@ 3.2 Experiments Results For Two Stage Algorithms
We experimentally validate the above theory, showing that a collective can be more successful against JTT and LfF compared to ERM on a synthetic 2D dataset and CIFAR-10 (Krizhevsky, 2009). These algorithms are explained in Definition 3 and 4 and the synthetic 2D dataset comprises points sampled i.i.d. from a uniform distribution on $\left(-1,1\right)^{2}$, labeled by the sign of their first coordinate $x_{1}$ (Figure 1(a)). The collective wants the points with a negative $x_{2}$ to be labeled $y^{*}$ and uses the strategy $\left\{\left(x_{1},x_{2}\right),y\right\}{\rightarrow}\left\{\left(x_{1},-\left|x_{2}\right|\right),y^{*}\right\}$. We also consider the multi-class classification problem of CIFAR-10. The collective transformation $g$, in the pixel space of integer values from 0 to 255, adds a perturbation of magnitude 2 to the value of every second pixel in every second row (Figure 1(a)). This transform is virtually invisible to the human eye. Technical details can be found in Appendix C.
The results on the 2D dataset (Figure 1(b)) and on CIFAR-10 (Figure 1(c)) show that for small $\alpha$, a collective achieves higher success in JTT and LfF than with the same $\alpha$ in ERM. Proposition 3 predicts this behavior for JTT. When $\alpha$ is small, the data bias causes ERM to misclassify collective members. This inaccuracy increases the collective population in the error set of first phase. This leads to a higher $P_{E\mid C}$ relatively to $P_{E}$ (Equation 14), which consequently leads to $\alpha_{\text{eff}}>\alpha$. This is no longer the case when $\alpha$ is large enough. As $\alpha$ rises, the accuracy of ERM on the collective increases, and the collective membership in the error set decreases. As a result, $P_{E\mid C}<P_{E}$ and $\alpha_{\text{eff}}<\alpha$, lowering the success. This effect starts at $\alpha\approx 0.2$ on the 2D dataset and at $\alpha\approx 0.1$ on CIFAR-10.
@@ 3.3 Iterative Re-weighting Algorithms and Collective Action on Validation Sets
We now turn to analyse algorithms that iteratively re-weight, focusing on CVaR-DRO. For these algorithms, the effective distribution at each step is chosen adversarially from the uncertainty set of distributions as discussed in Section 2.2. We explain the CVaR-DRO algorithm in definition 5. This causes the algorithm to cycle between fitting different parts of the distribution, terminating only when a high enough accuracy is reached on some validation set. As the validation set plays an important role in deciding when the algorithm terminates, we consider how varying $\alpha$ in the validation set can affect the collective success.
@@
In order to theoretically analyse the effect of the validation set on the final success of the collective, we look at an idealised version of a iterative re-weighting DRO algorithm. This ideal algorithm computes a sequence of classifiers ${\mathcal{F}}$, where each classifier $f^{\left(i\right)}$ is a Bayes optimal classifier for a distribution on which the previous classifier $f^{\left(i-1\right)}$ has the maximum error, starting with ${\mathcal{P}}_{0}$. The algorithm outputs the classifier $f{\in}{\mathcal{F}}$ that has the highest accuracy on the validation distribution affected by collective action, which is given by ${\mathcal{P}}_{V}=\beta P^{*}+(1-\beta){\mathcal{P}}_{0}$, where $\beta$ is the collective size in the validation set. A precise form of this algorithm is given in Algorithm 3.
Proposition 4.
@@
Figure 2(a) demonstrates the oscillating success after every epoch when training CVaR-DRO on the Waterbirds dataset Sagawa et al. (2020) with $\alpha{=}0.3$. These oscillations are due to the search over the distributions space ${\mathcal{Q}}_{{\mathcal{P}}_{\alpha}}$(Equation 6). Proposition 2 predicts that this space contains a distribution where the collective is stronger than in the observed distribution. The peaks of the oscillations suggests that CVaR-DRO was able to find such a distribution. The frequency and amplitude of the oscillations show that the resulting success is sensitive to when training stops, varying from minimal to maximal success.
@@
If the collective is able to affect the validation set, then increasing its proportion in the validation set will increase success, as suggested in Proposition 4. We show this experimentally by applying collective action in both training and validation set of the CIFAR-10 and Waterbirds datasets. Figure 3(b) and 3(d) show that when CVaR-DRO is used, the amount of collective in the validation set has a very large impact on the collective success compared to the collective size in the training set. In contrast, Figure 3(a) and 3(c) show that success in ERM is more sensitive to the collective size in the training set $\alpha$ rather than to the size in the validation set. With $\alpha=0.1$ in ERM, the collective achieves almost full success, while when $\alpha=0.001$, even with full control of the validations set, the success does not go over $0.4$.
@@ 4.1 Theoretical Results
(a) Collective signal
(b) On Class Automobile
(c) On Class Truck
@@ 4.2 Experimental Results With Simplicity Bias
(a) LMS-6 dataset
(b) Success on LMS-$k$
@@
Our results in Figure 4(b) and 4(c) show that as the correlation decreases, the success increases. The underlying intuition is that with higher correlation the simpler MNIST part alone can yield high training performance. As a result, SGD tends to ignore the CIFAR part, where the signal acts. In contrast, a weaker correlation between the MNIST part and the label necessitates the algorithm’s reliance on the CIFAR part, leading it to also learn the collective signal embedded therein. Note that for small $\alpha$ in Figure 4(b), the success with non-zero correlation is low because the collective is labeling automobiles as $y^{*}{=}\text{truck}$, resulting in competing signals.
@@ B.2 Proofs for results in Section 3
Then we restate and prove Proposition 4. See 4
@@ B.4 Example for theorem 5
Base distribution Building on the intuition from the MNIST-CIFAR example, we assume a distribution ${\mathcal{P}}_{0}(x_{1},x_{2},y)$ over $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{Y}$ where $\mathcal{X}_{i}$ is a set of size 10 and $\mathcal{Y}=\{+,y_{-}\}$. In the base distribution ${\mathcal{P}}_{0}$ we assume that both $x_{1},x_{2}$ are perfectly correlated with $y$ and so either is enough to determine the outcome. This can be seen as similar to MNIST-CIFAR where $x_{1}$ and $x_{2}$ correspond to the MNIST and CIFAR image respectively. We will also take the following:
@@
    $P_{1+}\left(x_{1}\right)$ is a uniform discrete probability over the set $X_{1+}$ which contains 5 values that correspond to the label $+$, i.e. $P_{1+}\left(x_{1}\in X_{1+}\right)=\frac{1}{5},P_{1}\left(x_{1}\notin X_{1+}\right)=0$.
@@
    $P_{1-}\left(x_{1}\right)$ is a uniform discrete probability over the set $X_{1-}$ that correspond to the label --.
@@
    $P_{2+}\left(x_{2}\right)$ is a uniform discrete probability over the set $X_{2+}$ that correspond to the label $+$.
@@
    $P_{2-}\left(x_{2}\right)$ is a uniform discrete probability over the set $X_{2-}$ that correspond to the label --.

### v2

@@ Abstract
Collective action in machine learning is the study of the control that a coordinated group can have over machine learning algorithms. While previous research has concentrated on assessing the impact of collectives against Bayes (sub)-optimal classifiers, this perspective is limited in that it does not account for the choice of learning algorithm. Classifiers seldom behave like Bayes classifiers and are influenced by the choice of learning algorithms along with their inherent biases. In this work, we initiate the study of how the choice of the learning algorithm plays a role in the success of a collective in practical settings. Specifically, we focus on distributionally robust optimization (DRO), popular for improving a worst group error, and on the ubiquitous stochastic gradient descent (SGD), due to its inductive bias for “simpler” functions. Our empirical results, supported by a theoretical foundation, show that the effective size and success of the collective are highly dependent on properties of the learning algorithm. This highlights the necessity of taking the learning algorithm into account when studying the impact of collective action in machine learning.
[…]
@@ 1 Introduction
Hardt et al. (2023) initiated the formal study into the success of different collective action strategies. Their work provided theoretical analysis based on the fractional size of the collective and on the properties of the signal with regards to the original data distribution. However, their results only hold for Bayes (sub)-optimal classifiers which do not immediately adapt to peculiarities of practical learning algorithms. In this work, we investigate two types of commonly used learning algorithms that exhibit distinct properties: (1) Distributional robustness and (2) Simplicity bias. We show how these properties lead to significantly different levels of collective success that are unexplained by prior work.
When the data is composed of multiple sub-populations, algorithms that optimise for average performance often perform poorly in minority sub-populations (Meinshausen and Bühlmann, 2015). A fairness-focused firm will want to ensure that their trained learning model performs well uniformly on all sub-populations, as opposed to being on-average good. Distributionally Robust Optimisation (DRO) is a family of algorithms designed to maximize this per-group accuracy (Hashimoto et al., 2018; Wang et al., 2020). We show that as a consequence, a small collective achieves higher success when the training algorithm performs DRO, compared to standard empirical risk minimization (ERM). Conversely, a large collective in the same settings achieves lower success, contradicting the expectation set by previous work regarding the effectiveness of large collectives.
Second, most machine learning algorithms used today are based on some form of gradient descent (GD). Such algorithms, including the popular Stochastic Gradient Descent (SGD), Adam (Kingma and Ba, 2015), and RMSProp, exhibit a preference for learning functions that are “simpler". This inductive bias of GD algorithms is popularly referred to as simplicity bias (Kalimeris et al., 2019; Shah et al., 2020). In practice, this preference results in the model “overlooking” certain complex features. We demonstrate that these overlooked features can be leveraged by a collective to design a strategy that will gain higher success compared to what is possible on a Bayes optimal classifier.
@@ 2.1 Collective Action
While there are several goals in the domain of algorithmic collective action defined in Hardt et al. (2023), in this work we will focus on the goal of collectively planting a signal with a features-label strategy. We believe that other important goals, like erasing signals, can also benefit from the techniques and observations of this paper, but we leave the detailed study of such settings to future work.
@@
The collective modifies their own data by planting a signal $\left(x,y\right){\rightarrow}\left(g\left(x\right){,}y^{*}\right)$${\mathcal{P}}^{*}$ defines the distribution of $\left({g\left({x}\right),y^{*}}\right)$ where $x\sim{\mathcal{P}}_{0}$. This has been defined as the feature-label strategy in prior work. The signal $g$ also defines the signal set ${\mathcal{X}}^{*}=\left\{g\left(x\right)|x\in{\mathcal{X}}\right\}$. For any distribution ${\mathcal{P}}$ over $X{\times}Y$, the Bayes optimal classifier on ${\mathcal{P}}$, which we denote as $f_{\mathcal{P}}$, is defined as
@@
Using the above four properties, they derive the following lower bound on the success of the collective.
@@ 2.2 Distributionally Robust Optimisation
where ${\mathcal{Q}}_{p}$ is an uncertainty set of distributions close to $p$ over which we want to control the risk, $\ell$ is the loss function, and $g_{\theta}$ is a function with parameters $\theta$. There are many possible definitions of the uncertainty set that the algorithm is controlling for. One possible choice is an $f$-divergence111 Definition 2 in the appendix. ball (Ali and Silvey, 1966; Csiszár, 1967).
@@ 3 Effective Size and Validation Control
The most intuitive parameter to predict the success of collective action is the fractional size of the collective $\alpha$; a larger collective will attain greater success. However, DRO algorithms assign different weights to different samples, rendering $\alpha$ inappropriate for predicting success. Instead, we introduce a correction to the collective size under a weighted distribution that we denote the effective size $\alpha_{\text{eff}}$, and show that DRO algorithms can yield $\alpha_{\text{eff}}>\alpha$.
@@ 3.1 Effective Collective Size
Note that if $w\left({x}\right){=}1$ for all $x$, then $\alpha_{\text{eff}}{=}\alpha$. Now, since $\alpha_{\text{eff}}$ is the collective size under the weighted distribution, then bounding the success for this algorithm, akin to Theorem 1, requires adding a corrective term $c$ for the non-weighted distribution, giving
@@
where $\Delta_{x}^{w}=\max_{y\in Y}\left({\mathcal{P}}^{\left(w\right)}\left(y\mid x\right)-{\mathcal{P}}^{\left(w\right)}\left(y^{*}\mid x\right)\right)$ and $\Delta_{x}$ is defined the same but for ${\mathcal{P}}_{0}$. Formal proof and definitions are in the appendix under Proposition B.3. Under certain circumstances we can have $c\leq 0$, implying a collective success greater than that guaranteed by Theorem 1 for $\alpha=\alpha_{\text{eff}}$. We provide such an example where $c\leq 0$ and $\alpha_{\text{eff}}\geq\alpha$ in Appendix B.1.
We now provide theoretical results, showing that DRO algorithms can increase the effective collective size $\alpha_{\text{eff}}$. Firstly, for any algorithm that targets a DRO defined by an $f$-divergence, we can say the following:
Proposition 3.1.
@@
Figure 2: Image transformation used by the collective. The effect of the signal is exaggerated for visualisation purposes and in practice it is invisible to the human eye.
@@ 3.2 Experiments Results for Two Stage Algorithms
We experimentally validate the above theory, showing that a collective can be more successful against JTT and LfF compared to ERM on a synthetic 2D dataset and CIFAR-10 (Krizhevsky, 2009). These algorithms are explained in Definition 3 and 4 and the synthetic 2D dataset comprises points sampled i.i.d. from a uniform distribution on $\left(-1,1\right)^{2}$, labeled by the sign of their first coordinate $x_{1}$ (Figure 1(a)). The collective wants the points with a negative $x_{2}$ to be labeled $y^{*}$ and uses the strategy $\left\{\left(x_{1},x_{2}\right),y\right\}{\rightarrow}\left\{\left(x_{1},-\left|x_{2}\right|\right),y^{*}\right\}$. We also consider the multi-class classification problem of CIFAR-10. The collective transformation $g$, in the pixel space of integer values from 0 to 255, adds a perturbation of magnitude 2 to the value of every second pixel in every second row (Figure 2). This transform is virtually invisible to the human eye. Technical details can be found in Appendix C.
The results on the 2D dataset (Figure 1(b)) and on CIFAR-10 (Figure 1(c)) show that for small $\alpha$, a collective achieves higher success in JTT and LfF than with the same $\alpha$ in ERM. Proposition 3.2 predicts this behavior for JTT. When $\alpha$ is small, the data bias causes ERM to misclassify collective members. This inaccuracy increases the collective population in the error set of first phase. This leads to a higher $P_{E\mid C}$ relatively to $P_{E}$ (Equation 14), which consequently leads to $\alpha_{\text{eff}}>\alpha$. This is no longer the case when $\alpha$ is large enough. As $\alpha$ rises, the accuracy of ERM on the collective increases, and the collective membership in the error set decreases. As a result, $P_{E\mid C}<P_{E}$ and $\alpha_{\text{eff}}<\alpha$, lowering the success. This effect starts at $\alpha\approx 0.2$ on the 2D dataset and at $\alpha\approx 0.1$ on CIFAR-10.
@@ 3.3 Iterative Re-weighting Algorithms and Collective Action on Validation Sets
We now turn to analyse algorithms that iteratively re-weight, focusing on CVaR-DRO. For these algorithms, the effective distribution at each step is chosen adversarially from the uncertainty set of distributions as discussed in Section 2.2. We explain the CVaR-DRO algorithm in definition 5 in the appendix. This causes the algorithm to cycle between fitting different parts of the distribution, terminating only when a high enough accuracy is reached on some validation set. As the validation set plays an important role in deciding when the algorithm terminates, we consider how varying $\alpha$ in the validation set can affect the collective success.
@@
In order to theoretically analyse the effect of the validation set on the final success of the collective, we look at an idealised version of a iterative re-weighting DRO algorithm. This ideal algorithm computes a sequence of classifiers ${\mathcal{F}}$, where each classifier $f^{\left(i\right)}$ is a Bayes optimal classifier for a distribution on which the previous classifier $f^{\left(i-1\right)}$ has the maximum error, starting with ${\mathcal{P}}_{0}$. The algorithm outputs the classifier $f{\in}{\mathcal{F}}$ that has the highest accuracy on the validation distribution affected by collective action, which is given by ${\mathcal{P}}_{V}=\beta P^{*}+(1-\beta){\mathcal{P}}_{0}$, where $\beta$ is the collective size in the validation set. A precise form of this algorithm is given in Algorithm 3 in the appendix.
Proposition 3.3.
@@
Figure 3(a) demonstrates the oscillating success after every epoch when training CVaR-DRO on the Waterbirds dataset Sagawa et al. (2020) with $\alpha{=}0.3$. These oscillations are due to the search over the distributions space ${\mathcal{Q}}_{{\mathcal{P}}_{\alpha}}$(Equation 6). Proposition 3.1 predicts that this space contains a distribution where the collective is stronger than in the observed distribution. The peaks of the oscillations suggests that CVaR-DRO was able to find such a distribution. The frequency and amplitude of the oscillations show that the resulting success is sensitive to when training stops, varying from minimal to maximal success.
@@
If the collective is able to affect the validation set, then increasing its proportion in the validation set will increase success, as suggested in Proposition 3.3. We show this experimentally by applying collective action in both training and validation set of the CIFAR-10 and Waterbirds datasets. Figure 4(b) and 4(d) show that when CVaR-DRO is used, the amount of collective in the validation set has a very large impact on the collective success compared to the collective size in the training set. In contrast, Figure 4(a) and 4(c) show that success in ERM is more sensitive to the collective size in the training set $\alpha$ rather than to the size in the validation set. With $\alpha=0.1$ in ERM, the collective achieves almost full success, while when $\alpha=0.001$, even with full control of the validations set, the success does not go over $0.4$.
@@ 4.1 Theoretical Results
(a) LMS-6 dataset
(b) Success on LMS-$k$
@@ 4.2 Experimental Results With Simplicity Bias
(a) Collective signal
(b) On Class Automobile
(c) On Class Truck
@@
Our results in Figure 6(b) and 6(c) show that as the correlation decreases, the success increases. The underlying intuition is that with higher correlation the simpler MNIST part alone can yield high training performance. As a result, SGD tends to ignore the CIFAR part, where the signal acts. In contrast, a weaker correlation between the MNIST part and the label necessitates the algorithm’s reliance on the CIFAR part, leading it to also learn the collective signal embedded therein. Note that for small $\alpha$ in Figure 6(b), the success with non-zero correlation is low because the collective is labeling automobiles as $y^{*}{=}\text{truck}$, resulting in competing signals.
@@ B.2 Proofs for Results in Section 3
Then we restate and prove Proposition 3.3. See 3.3
@@ B.4 Example for Theorem 2
Base distribution Building on the intuition from the MNIST-CIFAR example, we assume a distribution ${\mathcal{P}}_{0}(x_{1},x_{2},y)$ over $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{Y}$ where $\mathcal{X}_{i}$ is a set of size 10 and $\mathcal{Y}=\{y_{+},y_{-}\}$. In the base distribution ${\mathcal{P}}_{0}$ we assume that both $x_{1},x_{2}$ are perfectly correlated with $y$ and so either is enough to determine the outcome. This can be seen as similar to MNIST-CIFAR where $x_{1}$ and $x_{2}$ correspond to the MNIST and CIFAR image respectively. We will also take the following:
@@
    $P_{1+}\left(x_{1}\right)$ is a uniform discrete probability over the set $X_{1+}$ which contains 5 values that correspond to the label $y_{+}$, i.e. $P_{1+}\left(x_{1}\in X_{1+}\right)=\frac{1}{5},P_{1}\left(x_{1}\notin X_{1+}\right)=0$.
@@
    $P_{1-}\left(x_{1}\right)$ is a uniform discrete probability over the set $X_{1-}$ that correspond to the label $y_{-}$.
@@
    $P_{2+}\left(x_{2}\right)$ is a uniform discrete probability over the set $X_{2+}$ that correspond to the label $y_{+}$.
@@
    $P_{2-}\left(x_{2}\right)$ is a uniform discrete probability over the set $X_{2-}$ that correspond to the label $y_{-}$.